# A demonstration that classical gravity does not produce entanglement

### v1

@@ 1 Introduction
In [aziz2025classical], the authors dispute Claim. They purport to show that classical gravity can produce entanglement by providing a particular Hamiltonian, which they interpret as generating quantum communication and therefore violating an underlying LOCC assumption familiar in quantum information theory (because it violates the CC part: classical communication). Likewise, in [lin2025newtonian], the authors purport to show that classical gravity can produce entanglement by providing a particular Hamiltonian, which they interpret as generating non-local effects and therefore violating the same LOCC assumption (now because it violates the LO part: local operations). In fact, the latter invokes what we explicate in [huggett2023quantum] as a naive “Newtonian model” of GIE: an idealization in which the experiment is described as a bipartite system, with a direct, non-local interaction term, not a tripartite system that includes a gravitational mediator. So such a model is simply silent on any remark, like Claim, which assumes a gravitational mediator.
[aziz2025classical] (and [lin2025newtonian]) proceed by interpreting interaction Hamiltonians, and they accept a standard idealization that Newtonian gravity suffices to capture the role of gravity in the experiment protocol, rather than relativistic gravity. Objections to the original authors’ analysis of the Hamiltonian they provide have already appeared [marletto2025classical, marlettoETC2025classical, diosi2025no, di2025simple]. But classical gravity cannot produce entanglement on independent grounds, having nothing to do with a Hamiltonian formalism. In this note, we reiterate our demonstration that in the agreed regime of the experiment, if classical gravity is taken to be the Newtonian limit of classical Einstein gravity, then gravitational mediation simply will not induce entanglement. Claim is on solid grounds.
@@ 2 A Newton-Cartan Analysis
As further evidence for (and explication of) the claim that NCG captures the Newtonian limit of classical gravitational mediation, note that NCG can be quantized into a gauge field theory, whose quanta (“Newton-Cartan gravitons”) are infinite speed longitudinal bosons [christian1997exactly]. These, like the more familiar gravitons of quantized linear GR, form a third subsystem mediating gravity between a pair of massive subsystems. As we pointed out in [huggett2023quantum] this tripartite subsystem picture is essential for sensibly invoking LOCC, so must be respected. The operative sense of ‘local’ is that there is no direct interaction between the masses, but only with the mediator (indeed, our original point was that NCG shows that instantaneous propagation in a classical setting is not sufficient for mediating entanglement). That such a tripartite model arises from quantizing NCG is evidence that one can also safely take classical NCG to capture the Newtonian limit in the most perspicuous way for investigating classical mediation.555[dibiagio2025circuitlocalityrelativisticlocality, di2025simple] argue that LOCC (and other no-go theorems) require a stricter sense of ‘mediator’, which excludes gauge theories, including general relativity, and potentially NCG. We do not claim that our definition of ‘mediator’ is sufficient for LOCC (though it is necessary), only that it is sufficient for a more general notion of mediation, in line with what is expressed by the impressionistic equation 2 in [di2025simple]. Indeed, we take the discussion that closes [dibiagio2025circuitlocalityrelativisticlocality, §F] as tacit acknowledgment that the more general notion is (sometimes) what is physically relevant.

### v2

@@ 1 Introduction
In [aziz2025classical], the authors dispute Claim. They purport to show that classical gravity can produce entanglement by providing a particular Hamiltonian, which they interpret as generating quantum communication and therefore violating an underlying LOCC assumption familiar in quantum information theory (because it violates the CC part: classical communication). Likewise, in the original version (v.1) of [lin2025newtonian], the authors purported to show that classical gravity can produce entanglement by providing a particular Hamiltonian, which they interpret as generating non-local effects and therefore violating the same LOCC assumption (now because it violates the LO part: local operations). In fact, the latter invokes what we explicate in [huggett2023quantum] as a naive “Newtonian model” of GIE: an idealization in which the experiment is described as a bipartite system, with a direct, non-local interaction term, not a tripartite system that includes a gravitational mediator. So such a model is simply silent on any remark, like Claim, which assumes a gravitational mediator. [^5] After posting the initial version of this note to the ArXiv, we learned that shortly prior, a subsequent version of [lin2025newtonian] had reversed the authors’ original pronouncement. In the later version, the authors provide demonstrations in alignment with the standard reasoning that holds Newtonian gravity to be a mediator: “We show that the mesoscopic quantum bodies with their parities of mass quadrupoles interacting via Newtonian gravity can get entangled only if the parity of the gravitational tidal field sourced by the quantum body is also quantized” (v.2 abstract).
[aziz2025classical] (and [lin2025newtonian, v.1]) proceed by interpreting interaction Hamiltonians, and they accept a standard idealization that Newtonian gravity suffices to capture the role of gravity in the experiment protocol, rather than relativistic gravity. Objections to the original authors’ analysis of the Hamiltonian they provide have already appeared [marletto2025classical, marlettoETC2025classical, diosi2025no, di2025simple]. But classical gravity cannot produce entanglement on independent grounds, having nothing to do with a Hamiltonian formalism. In this note, we reiterate our demonstration that in the agreed regime of the experiment, if classical gravity is taken to be the Newtonian limit of classical Einstein gravity, then gravitational mediation simply will not induce entanglement. Claim is on solid grounds.
@@ 2 A Newton-Cartan Analysis
As further evidence for (and explication of) the claim that NCG captures the Newtonian limit of classical gravitational mediation, note that NCG can be quantized into a gauge field theory, whose quanta (“Newton-Cartan gravitons”) are infinite speed longitudinal bosons [christian1997exactly]. These, like the more familiar gravitons of quantized linear GR, form a third subsystem mediating gravity between a pair of massive subsystems. As we pointed out in [huggett2023quantum] this tripartite subsystem picture is essential for sensibly invoking LOCC, so must be respected. The operative sense of ‘local’ is that there is no direct interaction between the masses, but only with the mediator (indeed, our original point was that NCG shows that instantaneous propagation in a classical setting is not sufficient for mediating entanglement). That such a tripartite model arises from quantizing NCG is evidence that one can also safely take classical NCG to capture the Newtonian limit in the most perspicuous way for investigating classical mediation.666[dibiagio2025circuitlocalityrelativisticlocality, di2025simple] argue that LOCC (and other no-go theorems) require a stricter sense of ‘mediator’, which excludes gauge theories, including general relativity, and potentially NCG. We do not claim that our definition of ‘mediator’ is sufficient for LOCC (though it is necessary), only that it is sufficient for a more general notion of mediation, in line with what is expressed by the impressionistic equation 2 in [di2025simple]. Indeed, we take the discussion that closes [dibiagio2025circuitlocalityrelativisticlocality, §F] as tacit acknowledgment that the more general notion is (sometimes) what is physically relevant.